# The tidal evolution of anisotropic subhaloes: A new pathway to creating isotropic and cored satellites

### v1

@@ Abstract
pubyear: 2024pagerange: The tidal evolution of anisotropic subhaloes: A new pathway to creating isotropic and cored satellitesReferences
@@ 1 Introduction
The abundance and demographics of dark matter substructure is a sensitive probe of the nature of dark matter. Numerous ongoing efforts therefore focus on constraining the amount of dark matter substructure using a variety of different techniques, including annihilation and decay signals of dark matter particles (e.g., Strigari et al., 2007; Pieri et al., 2008; Ando et al., 2019), gravitational lensing (e.g., Vegetti et al., 2014; Hezaveh et al., 2016; Meneghetti et al., 2020), gaps in stellar streams (e.g., Carlberg, 2012; Erkal et al., 2016; Bonaca et al., 2019), and the abundance of satellite galaxies (e.g., Anderhalden et al., 2013; Nadler et al., 2021)
@@
This has prompted several studies based on high-resolution, idealised simulations to study the tidal evolution of subhaloes in a highly controlled setting that can avoid issues related to artificial disruption (Ogiya et al., 2019; Green & van den Bosch, 2019; Errani & Peñarrubia, 2020; Benson & Du, 2022; Errani & Navarro, 2021). One of the key insights from these studies, corroborated by analytical studies, is that the inner cusp of an NFW subhalo is always preserved during tidal evolution, and physical tidal disruption can never occur (van den Bosch et al., 2018; Errani & Peñarrubia, 2020; Errani & Navarro, 2021; Amorisco, 2021; Stücker et al., 2023). Another key result is that subhaloes evolve along ‘tidal tracks’ [^1] Actually, this insight goes back to Hayashi et al. (2003) and Peñarrubia et al. (2008), tight correlations between the bound mass fraction and parameters that characterise subhalo structure, such as peak circular velocity. This empirical fact implies that the density profile of a tidal remnant is fully specified by its initial (i.e., pre-infall) density profile and the fraction of its mass lost; it does not depend on how or when that mass was lost. Hence, all subhaloes with the same initial density profile evolve along the same tidal track, irrespective of the detailed host potential or the orbital parameters (Peñarrubia et al., 2010). idealised simulations have been used to obtain empirical fitting functions for the tidal tracks, which have been independently corroborated by several studies (e.g. Green & van den Bosch, 2019; Errani & Navarro, 2021; Amorisco, 2021; Benson & Du, 2022; Stücker et al., 2023) and have found wide applications in semi-analytic modelling of subhhaloes and satellite galaxies (e.g. Yang et al., 2020; Jiang et al., 2021; Ahvazi et al., 2023; Green et al., 2022; Folsom et al., 2023; Du et al., 2024). Furthermore, it has been suggested that, akin to the energy-lowering isothermal King model (King, 1966), the process of subhalo mass stripping can be accurately approximated as consecutive ‘shell peeling’ in the energy space (e.g. Drakos et al., 2017, 2020). As the precise ‘energy truncation’ is uniquely determined by the host halo properties and subhalo orbital parameters, the tidal mass loss rate can be analytically predicted by knowing only the subhalo initial density profiles without regard to its internal velocity-distribution structure (Drakos et al., 2022; Stücker et al., 2023).
@@
However, high-resolution cosmological simulations have shown that, in a $\Lambda$CDM cosmology, the velocity structure of dark matter haloes clearly deviates from isotropy (e.g., Diemand et al., 2007a; Navarro et al., 2010; Ludlow et al., 2011; Klypin et al., 2016; He et al., 2024) [^2] Haloes are also found to exhibit velocity anisotropy in alternative DM paradigms, including warm dark matter (Polisensky & Ricotti, 2015) and self-interacting dark matter (e.g. Chua et al., 2021; Shen et al., 2021).. These studies find that haloes overall are radially anisotropic. More specifally, the velocity anisotropy,
@@
relating the tangential and radial velocity dispersions profiles, $\sigma_{\rm t}(r)$ and $\sigma_{\rm r}(r)$, respectively, typically increases from $\beta\sim 0.2$ near the centre to $\beta\sim 0.3$ around the scale radius, where ${\rm d}\log\rho/{\rm d}\log r=-2$, and than falling again to $\sim 0.2$ near the virial radius. However, the halo-to-halo variance is large, especially in the halo outskirts, and there are some hints that the radial anisotropy is larger at higher redshifts (Klypin et al., 2016); there also appears to be clear halo-mass-dependence in $\beta(r)$ (He et al., 2024).
Figure 1: Initial density (top panel), velocity anisotropy (middle), and 3D velocity dispersion profiles (bottom) of all nine subhaloes discussed in the text (colour coded as indicated). The latter is normalised by $\sigma_{0}\equiv\sqrt{Gm_{\text{s,0}}/r_{\text{s,0}}}$. All subhaloes have identical, phase-space-truncated density profiles, but differ in their velocity anisotropy profiles. Note how subhaloes that are more radially anisotropic have a higher velocity dispersion in their central regions, and how the Osipkov-Merritt (OM) model transitions from being isotropic in the center to maximally radially anisotropic ($\beta=1.0$) near the halo’s truncation radius.
@@
Figure 2: Tidal evolution of an isotropic subhalo resolved with $N_{\text{par}}=10^{7}$ particles on a circular orbit with $\mathcal{R}_{E}=0.25$. Different columns correspond to different epochs, as indicated. Top row: Face-on density projection of all subhalo particles, colour-coded by surface density. Grey and green, dashed circles indicate the virial radius of the host halo and the circular orbit of the subhalo, respectively. Middle and Bottom rows: Face-on and edge-on zoom-in projections of the density of instantaneous bound (colour-coded) and unbound (black) particles. The blue, solid circle indicates the initial virial radius of the subhalo, anchored on its instantaneous center of mass (marked by a cross), and the green, dashed curve marks the subhalo orbit. The bound mass fraction, $f_{\rm bound}$, at each epoch is indicated in the panels in the middle row.
@@ 2.1 Initial conditions and subhalo orbits
Different subhaloes only differ in their velocity anisotropy profiles. Subhalo 1 has $\beta(r)=0$ and is thus isotropic throughout. Subhaloes 2 to 8 have a constant velocity anisotropy of $\beta(r)=-0.5,+0.1,+0.2,+0.25,+0.3,+0.4,$ and $+0.5$, respectively. Finally, subhalo 9 has an Osipkov–Merritt (OM) distribution function (Osipkov, 1979; Merritt, 1985) which implies that
@@
with $r_{\text{a}}$ the anisotropic radius. Hence, this system has an isotropic center ($r\ll r_{\text{a}}$) and becomes more and more radially anisotropic at larger radii. Unless specifically stated otherwise, we set $r_{\text{a}}=r_{\text{s,0}}$. Throughout this paper we refer to subhaloes 1, 2, 3–8, and 9 as the isotropic subhalo, the tangentially anisotropic subhalo, the radially anisotropic subhaloes, and the OM subhalo, respectively. Their (identical) density profiles, $\rho(r)$, anistropy profiles, $\beta(r)$, and 3D velocity dispersion profiles, $\sigma_{\rm 3D}(r)$ are shown in Fig. 1.
The ICs for each of these subhaloes are generated using the new, open-source Python code PIANISTpy (Particle Initial conditions with ANISoTropy) described in Chiang & van den Bosch (2024). This code, which we specifically developed for this project, generates equilibrium initial conditions for spherical, double-power-law density profiles that are truncated in phase-space similar to the energy truncation methodology of Drakos et al. (2017), and that can have a wide range of different anisotropy profiles. By design, the self-consistent truncation of the system’s distribution function does not alter the underlying velocity anisotropy profile $\beta(r)$, nor does it introduce any truncation-related instability, as for example in Weinberg (2022). Importantly, the ICs require no further relaxation as the particle orbit sampling is exact, free of inherent ‘imperfections’ that hamper approximate iterative schemes like the method of Schwarzschild (1979).
@@
In this study, we vary $\mathcal{R}_{E}$ over the range $[0.15,1.0]$ and mainly consider orbital eccentricities of $e=0$ (i.e., a circular orbit), $0.45$ and $0.9$. The orbital eccentricity is related to the orbital circularity, $\eta$, defined as the ratio of $L_{\rm orb}$, and the angular momentum $L_{\rm circ}(E_{\rm orb})$ corresponding to a circular orbit of energy $E_{\rm orb}$. For our host halo potential and $\mathcal{R}_{E}=0.25$, eccentricities of 00, $0.45$ and $0.9$ correspond to orbital circularities of $1.0$, $0.78$ and $0.17$, respectively.
@@ 2.2 Numerical simulations
All simulations presented in this work are performed using the code gamer-2 (Schive et al., 2018), which supports hybrid CPU/GPU parallelisation and adaptive mesh refinement (AMR) with octree data structure. We adopt a simulation box of dimensions $800r_{\text{s,0}}\times 800r_{\text{s,0}}\times 800r_{\text{s,0}}$, covered by $128^{3}$ root grids and nine refinement levels, achieving a maximum spatial resolution of $r_{\text{res}}=0.012r_{\text{s,0}}$. For all our simulations we adopt the standard AMR strategy that further refines a cell if the total DM particle number therein exceeds four [^3] Architecturally, gamer-2 groups $8\times 8\times 8$ cells as a patch and performs refinement on a patch-by-patch basis. In addition, refinement levels of adjacent patches can vary by at most one level. This design differs from fully-threaded-tree AMR codes such as RAMSES (Teyssier, 2002) that perform the refinement on a cell-by-cell basis. We have verified that our simulation results remain unchanged and numerically convergent under alternative mass+geometry-based AMR strategies.. The large box size ensures that tidally stripped particles always remain sufficiently distant from simulation boundaries so as to minimise any possible numerical artifacts originating from the outflow boundary conditions.
At each integration time step, the subhalo’s instantaneous center of mass (CoM) position and velocity are computed on-the-fly from the $5\%$ most bound particles, following the iterative procedure outlined in Section 2.3 of van den Bosch & Ogiya (2018). In essence, during each iteration the subhalo self-potential is first computed from the instantaneous bound DM particles. Next, each individual DM particle’s total energy, defined with respect to the subhalo’s CoM, is updated and sorted, after which the new CoM position and velocity are determined using the $5\%$ most bound particles. This process is iterated until the changes in the CoM position and velocity are smaller than $10^{-4}r_{\rm vir}$ and $10^{-4}V_{\text{vir}}$, respectively. The bound mass fraction is defined as
@@
We have carefully verified that the adopted numerical scheme is conservatively adequate to resolve the dynamical evolution of (initially) cuspy subhaloes and guarantees numerical convergence of the tidal evolution of our subhaloes down to $f_{\rm bound}\leq 10^{-5}$. Additional numerical tests on subhalo stability, between gamer-2 and treecode, and against different initial condition setups are presented in Appendix A.
@@ 3.1 Tidal stripping of anisotropic subhaloes
In order to probe the impact of internal velocity anisotropy on the tidal mass loss rate, we evolve all 9 subhaloes, which only differ in their initial velocity anisotropy, on circular orbits with $\mathcal{R}_{E}=0.25$, along which the tidal field strength remains time-invariant. Figs. 2 and 3 show snapshots of the tidal evolution of the isotropic subhalo and the radially anisotropic subhalo with $\beta=0.25$, respectively. The top rows show the face-on (defined with respect to the orbital plane of the subhalo) density projection of all subhalo particles, colour-coded based on the total projected particle density. Green and gray dashed circles indicate the subhalo orbit and the virial radius of the host halo, respectively. The middle and bottom rows of panels show face-on and edge-on zoom-ins on the subhalo. Particles that are bound are colour-coded by projected density, while particles that are (instantaneously) unbound are black. The blue circle indicates the virial radius of the initial subhalo, centered on the instantaneous CoM of the subhalo, and is shown for comparison.
@@
This is further accentuated in Fig. 4, which compares the evolution of the bound mass fraction, $f_{\rm bound}(t)$, for all nine subhaloes evolved along two different orbits with $\mathcal{R}_{E}=0.25$: a circular orbit with $e=0$ (top panel) and a highly eccentric orbit with $e=0.9$ (bottom panel). As is evident, the amount of mass loss experienced by the subhaloes increases with increasing $\beta$; while the tangentially anisotropic subhalo with $\beta=-0.5$ experiences less mass loss than its isotropic counterpart, the radially anisotropic subhaloes experience an enhanced mass loss. By the end of a single orbit (indicated by the vertical, light-gray dotted line), the bound mass fractions of the various subhaloes already differ by up to a factor of $\sim 2$ compared to the standard ‘benchmark’ isotropic subhalo. With the exception of the OM subhalo, this difference increases monotonically with time, and after a Hubble time (vertical dark-gray dashed line), the bound mass fraction can be different by orders of magnitude, especially along the eccentric orbit. In fact, the subhaloes with $\beta=0.5$ completely disrupt [^4] Operationally, we define disruption as the epoch when we can no longer identify a bound remnant, which typically happens shortly after the bound mass fraction drops below $10^{-5}$. within under 7 (5) Gyr in the case of the circular (eccentric) orbit. For comparison, the tangentially anisotropic subhalo still has a bound mass fraction of $\sim 0.05$ ($\sim 0.01$) after a Hubble time. In the case of the OM subhalo, the mass loss starts out even more extreme than for the $\beta=+0.5$ subhalo. However, after roughly one orbital period, when $f_{\rm bound}$ has dropped below $\sim 0.1$, the bound mass fraction of the OM subhalo starts to asymptote towards that of the isotropic subhalo. This behavior fits in with the overall trend, in that the outskirts of the OM subhalo are initially even more radially anisotropic than that of the constant $\beta=+0.5$ subhalo (see Fig. 1). However, once these outskirts have been stripped off, the remaining remnant is close to isotropic and its mass loss starts to resemble that of the isotropic subhalo.
@@
Figure 6: Fraction of constituent particle orbits, in the initial, unstripped subhalo, with instantaneous radius $r_{\text{par}}<r_{\rm t}$ that have an apo-center radius $r_{\text{apo}}>r_{\rm t}$ as a function of a fictitious tidal radius, $r_{\rm t}$. Curves are colour-coded by the subhalo initial anisotropy as indicated. As discussed in the text, this fraction is indicative of the mass that is likely to be stripped within roughly one dynamical time. Since more radial orbits have larger apocenters, subhaloes that are more radially anisotropic undergo more rapid stripping.
One can understand the strong dependence of the tidal mass loss on the internal velocity anisotropy by considering their orbital composition. Fig. 6 shows the fraction of particles, $f_{\rm p}$, with instantaneous radii $r_{\text{par}}<r_{\rm t}$ in the initial conditions that have apo-centric distances (computed using Eq. [6]) $r_{\rm apo}>r_{\rm t}$, as a function of some assumed value of the tidal radius, $r_{\rm t}$. Hence, these are the particles that are currently inside the tidal radius, but will orbit outside of $r_{\rm t}$, and thus be stripped, within less than one dynamical time. We can thus think of $f_{\rm p}$ as representative of the mass fraction that will be stripped off within about one dynamical time. Admittedly, this is crude, as it ignores the fact that prior stripping will have modified the density profile due to re-virialisation, and that $r_{\rm t}$ might well evolve in one dynamical time, but it suffices for the purpose of highlighting how the internal anisotropy impacts the stripping rate. Different colours correspond to the different initial anisotropies, as indicated. There is a clear trend of increasing $f_{\rm p}$ with increasing $\beta$. Particles in more radially anisotropic subhaloes are on more radial orbits, and thus, for a given orbital energy, have larger apo-centers, which in turn implies that they are more easily stripped. Note that this is a self-enhancing effect; once more mass is stripped off, the system re-virialises more strongly, which causes it to puff-up more strongly, which implies that the tidal radius moves in further, which promotes more stripping (see van den Bosch et al., 2018). This self-enhancement explains why the $f_{\rm bound}(t)$ curves for the various constant anisotropies in Fig. 4 diverge from each other as time progresses.

### v2

@@ Abstract
pubyear: 2025pagerange: The tidal evolution of anisotropic subhaloes: A new pathway to creating isotropic and cored satellitesReferences
@@ 1 Introduction
The abundance and demographics of dark matter substructure is a sensitive probe of the nature of dark matter. Numerous ongoing efforts therefore focus on constraining the amount of dark matter substructure using a variety of different techniques, including annihilation and decay signals of dark matter particles (e.g., Strigari et al., 2007; Pieri et al., 2008; Ando et al., 2019), gravitational lensing (e.g., Vegetti et al., 2014; Hezaveh et al., 2016; Meneghetti et al., 2020), gaps in stellar streams (e.g., Carlberg, 2012; Erkal et al., 2016; Bonaca et al., 2019), and the abundance of satellite galaxies (e.g., Anderhalden et al., 2013; Nadler et al., 2021).
@@
This has prompted several studies based on high-resolution, idealised simulations to study the tidal evolution of subhaloes in a highly controlled setting that can avoid issues related to artificial disruption (Ogiya et al., 2019; Green & van den Bosch, 2019; Errani & Peñarrubia, 2020; Benson & Du, 2022; Errani & Navarro, 2021). One of the key insights from these studies, corroborated by analytical studies, is that the inner cusp of an NFW subhalo is always preserved during tidal evolution, and physical tidal disruption can never occur (van den Bosch et al., 2018; Errani & Peñarrubia, 2020; Errani & Navarro, 2021; Amorisco, 2021; Stücker et al., 2023). Another key result is that subhaloes evolve along ‘tidal tracks’ [^1] Actually, this insight goes back to Hayashi et al. (2003) and Peñarrubia et al. (2008), tight correlations between the bound mass fraction and parameters that characterise subhalo structure, such as peak circular velocity. This empirical fact implies that the density profile of a tidal remnant is fully specified by its initial (i.e., pre-infall) density profile and the fraction of its mass lost; it does not depend on how or when that mass was lost. Hence, all subhaloes with the same initial density profile evolve along the same tidal track, irrespective of the detailed host potential or the orbital parameters (Peñarrubia et al., 2010). Idealised simulations have been used to obtain empirical fitting functions for the tidal tracks, which have been independently corroborated by several studies (e.g. Green & van den Bosch, 2019; Errani & Navarro, 2021; Amorisco, 2021; Benson & Du, 2022; Stücker et al., 2023) and have found wide applications in semi-analytic modelling of subhhaloes and satellite galaxies (e.g. Yang et al., 2020; Jiang et al., 2021; Ahvazi et al., 2023; Green et al., 2022; Du et al., 2024; Folsom et al., 2025). Furthermore, it has been suggested that, akin to the energy-lowering isothermal King model (King, 1966), the process of subhalo mass stripping can be accurately approximated as consecutive ‘shell peeling’ in energy space (e.g. Drakos et al., 2017, 2020; Errani et al., 2022). As the precise ‘energy truncation’ is uniquely determined by the host halo properties and subhalo orbital parameters, the tidal mass loss rate can be analytically predicted by knowing only the subhalo initial density profiles without regard to its internal velocity-distribution structure (Drakos et al., 2022; Stücker et al., 2023).
@@
[…]
@@
relating the tangential and radial velocity dispersion profiles, $\sigma_{\rm t}(r)$ and $\sigma_{\rm r}(r)$, respectively, typically increases from $\beta\sim 0.2$ near the centre to $\beta\sim 0.3$ around the scale radius, where ${\rm d}\log\rho/{\rm d}\log r=-2$, and than falling again to $\sim 0.2$ near the virial radius. However, the halo-to-halo variance is large, especially in the halo outskirts, and there are some hints that the radial anisotropy is larger at higher redshifts (Klypin et al., 2016); there also appears to be clear halo-mass-dependence in $\beta(r)$ (He et al., 2024).
Figure 1: Initial density (top panel), velocity anisotropy (middle), and 3D velocity dispersion profiles (bottom) of all nine subhaloes discussed in the text (colour coded as indicated). The latter is normalised by $\sigma_{0}\equiv\sqrt{Gm_{\text{s,0}}/r_{\text{s,0}}}$. All subhaloes have identical, phase-space-truncated density profiles, but differ in their velocity anisotropy profiles. Note how subhaloes that are more radially anisotropic have a higher velocity dispersion in their central regions, and how the Osipkov–Merritt (OM) model transitions from being isotropic in the centre to maximally radially anisotropic ($\beta=1.0$) near the halo’s truncation radius.
@@
Figure 2: Tidal evolution of an isotropic subhalo resolved with $N_{\text{par}}=10^{7}$ particles on a circular orbit with $\mathcal{R}_{E}=0.25$. Different columns correspond to different epochs, as indicated. Top row: Face-on density projection of all subhalo particles, colour-coded by surface density. Grey and green, dashed circles indicate the virial radius of the host halo and the circular orbit of the subhalo, respectively. Middle and Bottom rows: Face-on and edge-on zoom-in projections of the density of instantaneous bound (colour-coded) and unbound (black) particles. The blue, solid circle indicates the initial virial radius of the subhalo, anchored on its instantaneous centre of mass (marked by a cross), and the green, dashed curve marks the subhalo orbit. The bound mass fraction, $f_{\rm bound}$, at each epoch is indicated in the panels in the middle row.
@@ 2.1 Initial conditions and subhalo orbits
Different subhaloes only differ in their velocity anisotropy profiles. Subhalo 1 has $\beta(r)=0$ and is thus isotropic throughout. Subhaloes 2 to 8 have a constant velocity anisotropy of $\beta(r)=-0.5,+0.1,+0.2,+0.25,+0.3,+0.4,$ and +0.5+0.5, respectively. Finally, subhalo 9 has an Osipkov–Merritt (OM) distribution function (Osipkov, 1979; Merritt, 1985) which implies that
@@
with $r_{\text{a}}$ the anisotropic radius. Hence, this system has an isotropic centre ($r\ll r_{\text{a}}$) and becomes more and more radially anisotropic at larger radii. Throughout this paper, we set $r_{\text{a}}=1.4r_{\text{s,0}}$. We also refer to subhaloes 1, 2, 3–8, and 9 as the isotropic subhalo, the tangentially anisotropic subhalo, the radially anisotropic subhaloes, and the OM subhalo, respectively. Their (identical) density profiles, $\rho(r)$, anistropy profiles, $\beta(r)$, and 3D velocity dispersion profiles, $\sigma_{\rm 3D}(r)$ are shown in Fig. 1.
The ICs for each of these subhaloes are generated using the new, open-source Python code PIANISTpy (Particle Initial conditions with ANISoTropy) described in Chiang et al. (in preparation). This code, which we specifically developed for this project, generates equilibrium initial conditions for spherical, double-power-law density profiles that are truncated in phase-space similar to the energy truncation methodology of Drakos et al. (2017), and that can have a wide range of different anisotropy profiles. By design, the self-consistent truncation of the system’s distribution function does not alter the underlying velocity anisotropy profile $\beta(r)$, nor does it introduce any truncation-related instability, as for example in Weinberg (2023). Importantly, the ICs require no further relaxation as the particle orbit sampling is exact, free of inherent ‘imperfections’ that hamper approximate iterative schemes like the method of Schwarzschild (1979).
@@
In this study, we vary $\mathcal{R}_{E}$ over the range $[0.15,1.0]$ and mainly consider orbital eccentricities of $e=0$ (i.e., a circular orbit), $0.45$ and $0.9$. The orbital eccentricity is related to the orbital circularity, $\eta$, defined as the ratio of $L_{\rm orb}$, and the angular momentum $L_{\rm circ}(E_{\rm orb})$ corresponding to a circular orbit of energy $E_{\rm orb}$. For our host halo potential and $\mathcal{R}_{E}=0.25$, eccentricities of 0, $0.45$ and $0.9$ correspond to orbital circularities of $1.0$, $0.78$ and $0.17$, respectively.
@@ 2.2 Numerical simulations
All simulations presented in this work are performed using the code gamer-2 (Schive et al., 2018), which supports hybrid CPU/GPU parallelisation and adaptive mesh refinement (AMR) with octree data structure. We adopt a simulation box of dimensions $800r_{\text{s,0}}\times 800r_{\text{s,0}}\times 800r_{\text{s,0}}$, covered by $128^{3}$ root grids and nine refinement levels, achieving a maximum spatial resolution of $\Delta x_{\text{min}}=0.012r_{\text{s,0}}$. For all our simulations we adopt the standard AMR strategy that further refines a cell if the total dark matter particle number therein exceeds four [^3] Architecturally, gamer-2 groups $8\times 8\times 8$ cells as a patch and performs refinement on a patch-by-patch basis. In addition, refinement levels of adjacent patches can vary by at most one level. This design differs from fully-threaded-tree AMR codes such as RAMSES (Teyssier, 2002) that perform the refinement on a cell-by-cell basis. We have verified that our simulation results remain unchanged and numerically convergent under alternative mass+geometry-based AMR strategies.. The large box size ensures that tidally stripped particles always remain sufficiently distant from simulation boundaries so as to minimise any possible numerical artifacts originating from the outflow boundary conditions.
At each integration time step, the subhalo’s instantaneous centre of mass (CoM) position and velocity are computed on-the-fly from the $5\%$ most bound particles, following the iterative procedure outlined in Section 2.3 of van den Bosch & Ogiya (2018). In essence, during each iteration the subhalo self-potential is first computed from the instantaneous bound dark matter particles. Next, each individual dark matter particle’s total energy, defined with respect to the subhalo’s CoM, is updated and sorted, after which the new CoM position and velocity are determined using the $5\%$ most bound particles. This process is iterated until the changes in the CoM position and velocity are smaller than $10^{-4}r_{\rm vir}$ and $10^{-4}V_{\text{vir}}$, respectively. The bound mass fraction is defined as
@@
We have carefully verified that the adopted numerical scheme is conservatively adequate to resolve the dynamical evolution of (initially) cuspy subhaloes and guarantees numerical convergence of the tidal evolution of our subhaloes down to $f_{\rm bound}\leq 10^{-5}$. Additional numerical tests on subhalo stability, between gamer-2 and treecode, and against different initial condition setups are presented in Appendix A. Further comprehensive convergence tests can be found in Chiang et al. (2025).
@@ 3.1 Tidal stripping of anisotropic subhaloes
In order to probe the impact of internal velocity anisotropy on the tidal mass loss rate, we evolve all 9 subhaloes, which only differ in their initial velocity anisotropy, on circular orbits with $\mathcal{R}_{E}=0.25$, along which the tidal field strength remains time-invariant. Figs. 2 and 3 show snapshots of the tidal evolution of the isotropic subhalo and the radially anisotropic subhalo with $\beta=0.25$, respectively. The top rows show the face-on (defined with respect to the orbital plane of the subhalo) density projection of all subhalo particles, colour-coded based on the total projected particle density. Green and gray dashed circles indicate the subhalo orbit and the virial radius of the host halo, respectively. The middle and bottom rows of panels show face-on and edge-on zoom-ins on the subhalo. Particles that are bound are colour-coded by projected density, while particles that are (instantaneously) unbound are black. The blue circle indicates the virial radius of the initial subhalo, centred on the instantaneous CoM of the subhalo, and is shown for comparison. Note that the subhalo remnant drifts inward slightly over time. This is not due to dynamical friction against the host halo, which is absent since the latter is modeled as a static, analytical potential, but rather due to self-friction against the tidally stripped material (see Miller et al., 2020).
@@
This is further accentuated in Fig. 4, which compares the evolution of the bound mass fraction, $f_{\rm bound}(t)$, for all nine subhaloes evolved along two different orbits with $\mathcal{R}_{E}=0.25$: a circular orbit with $e=0$ (top panel) and a highly eccentric orbit with $e=0.9$ (bottom panel) [^4] The dips in the $f_{\rm bound}(t)$ curves in the lower panel occur near peri-centric passages, when some constituent particles are temporarily unbound from the subhalo by the strong and impulsive tidal shocks but then become bound again during subsequent re-virialisation away from the peri-centre (see van den Bosch & Ogiya, 2018).. As is evident, the amount of mass loss experienced by the subhaloes increases with increasing $\beta$; while the tangentially anisotropic subhalo with $\beta=-0.5$ experiences less mass loss than its isotropic counterpart, the radially anisotropic subhaloes experience an enhanced mass loss. By the end of a single orbit (indicated by the vertical, light-gray dotted line), the bound mass fractions of the various subhaloes already differ by up to a factor of $\sim 2$ compared to the standard ‘benchmark’ isotropic subhalo. With the exception of the OM subhalo, this difference increases monotonically with time, and after a Hubble time (vertical dark-gray dashed line), the bound mass fraction can be different by orders of magnitude, especially along the eccentric orbit. In fact, the subhaloes with $\beta=0.5$ completely disrupt [^5] Operationally, we define disruption as the epoch when we can no longer identify a bound remnant with $N_{\text{bound}}\leq 10$ (i.e. $f_{\rm bound}\leq 10^{-6}$). We have verified that our results are insensitive to the exact choice of disruption threshold, given that $f_{\rm bound}$ experiences a run-away decline after the onset of physical core disruption. within under 7 (5) Gyr in the case of the circular (eccentric) orbit. For comparison, the tangentially anisotropic subhalo still has a bound mass fraction of $\sim 0.05$ ($\sim 0.01$) after a Hubble time. In the case of the OM subhalo, the mass loss starts out even more extreme than for the $\beta=+0.5$ subhalo. However, after roughly one orbital period, when $f_{\rm bound}$ has dropped below $\sim 0.1$, the bound mass fraction of the OM subhalo starts to asymptote towards that of the isotropic subhalo. This behavior fits in with the overall trend, in that the outskirts of the OM subhalo are initially even more radially anisotropic than that of the constant $\beta=+0.5$ subhalo (see Fig. 1). However, once these outskirts have been stripped off, the remaining remnant is close to isotropic and its mass loss starts to resemble that of the isotropic subhalo.
@@
Figure 6: Fraction of constituent particle orbits, in the initial, unstripped subhalo, with instantaneous radius $r_{\text{par}}<r_{\rm t}$ that have an apo-centre radius $r_{\text{apo}}>r_{\rm t}$ as a function of a fictitious tidal radius, $r_{\rm t}$. Curves are colour-coded by the subhalo initial anisotropy as indicated. As discussed in the text, this fraction is indicative of the mass that is likely to be stripped within roughly one dynamical time. Since more radial orbits have larger apo-centres, subhaloes that are more radially anisotropic undergo more rapid stripping.
One can understand the strong dependence of the tidal mass loss on the internal velocity anisotropy by considering their orbital composition. Fig. 6 shows the fraction of particles, $f_{\rm p}$, with instantaneous radii $r_{\text{par}}<r_{\rm t}$ in the initial conditions that have apo-centric distances (computed using Eq. (6)) $r_{\rm apo}>r_{\rm t}$, as a function of some assumed value of the tidal radius, $r_{\rm t}$. Hence, these are the particles that are currently inside the tidal radius, but will orbit outside of $r_{\rm t}$, and thus be stripped, within less than one dynamical time. We can thus think of $f_{\rm p}$ as representative of the mass fraction that will be stripped off within about one dynamical time. Admittedly, this is crude, as it ignores the fact that prior stripping will have modified the density profile due to re-virialisation, and that $r_{\rm t}$ might well evolve in one dynamical time, but it suffices for the purpose of highlighting how the internal anisotropy impacts the stripping rate. Different colours correspond to the different initial anisotropies, as indicated. There is a clear trend of increasing $f_{\rm p}$ with increasing $\beta$. Particles in more radially anisotropic subhaloes are on more radial orbits, and thus, for a given orbital energy, have larger apo-centres, which in turn implies that they are more easily stripped. Note that this is a self-enhancing effect; once more mass is stripped off, the system re-virialises more strongly, which causes it to puff-up more strongly, which implies that the tidal radius moves in further, which promotes more stripping (see van den Bosch et al., 2018). This self-enhancement explains why the $f_{\rm bound}(t)$ curves for the various constant anisotropies in Fig. 4 diverge from each other as time progresses.